Questions Water is howing through a circular pipe with a radius of 0.30m at 340ms. If the pipes radius tapers down to 020m, show work please
Answer #1

Cross sectional area of a circular pipe having radius r meter is,

A = Πr2

Now, according to equation of continuity,

Av = constant.

Hare, v = velocity of the liquid.

Given in the question,

At radius r1 = 0.30 m , velocity , v1 = 3.40 m/s

Let, ar radius , r2 = 0.20 m, velocity is v2 .

Thus from equation of continuity,

r1​​​​​​2v1 = r22v2

Or, 0.30×0.30×3.40 = 0.20×0.20 × v2

Or, v2 = (0.30×0.30 × 3.40 )/0.20×0.20 = 5.1 m/s

Answer : 7.65 m/s